OpenStudy (anonymous):

FAN & MEDAL to who helps me solve: 2f(5f-2)-10(f^2-3f+6)=-8f(f+4)+4(2f^2-7f)

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

You have: \[2x(5x-2)-10(x^2-3x+6)=-8x(x+4)+4(2x^2-7x)\] Now we write the linear polynomial on the left hand side in standard form: \[26x-60=-8x(x+4)+4(2x^2-7x)\] Now we do the same on the right side: \[26x-60=-60x\] Add 60x +60 to both sides: \[86x=60\] Divide both sides with 86 \[x=30/43\]

OpenStudy (anonymous):

I used x instead of "f", this can simply be changed.

OpenStudy (anonymous):

where is 26 coming from @tommynaiter

OpenStudy (anonymous):

\[2f(5f-2)-10(f^2-3f+6)=-8f(f+4)+4(2f^2-7f)\]\[10f^2-4f-10f^2-30f-16=-8f^2-32+8f^2-28\]\[-34f-16=60\]

OpenStudy (anonymous):

Lets look at the left side. We have \[2x(5x-2)-10(x^2-3x+6)\] First i want to remove the brackets (). I just multiply in: \[10x^2 -4x-10(x^2-3x+6)=10x^2-4x-10x^2+30x-60=-4x+30x-60\] \[=26x-60\]

OpenStudy (anonymous):

this is what i keep getting

OpenStudy (anonymous):

You need to remember that it is -10. So in order to remove the negative, we have to change + to - and - to +, in the parentheses. \[-10(x^2-3x+6)=10(-x^2+3x-6)=-10x^2+30x-60\]

OpenStudy (anonymous):

i think im overthinking it because i dont get what you're saying

OpenStudy (anonymous):

Can you follow me here? \[2x(5x-2)=2x*5x-2x*2=10x^2-4x\] \[-10(x^2-3x+6)=-10*x^2+(-10)*(-3x)+6*(-10)=-10x^2+30x-60\]

OpenStudy (anonymous):

yeah, on the first line you distributed, but then idk what you did on the second line, i usually go through and distribute the whole line, then the next like i combine like terms and thats what im having trouble with, combining them and then everything after